# DeepDistill: Enhancing LLM Reasoning Capabilities via Large-Scale Difficulty-Graded Data Training

### v2

@@ 3.2 Data Selection
2:Category-specific verify_score threshold and fixed CV threshold (0.05)
@@
11:     if CV $>$ 0.05 then $\triangleright$ Challenging query
@@
      If the CV exceeds the threshold of 0.05, the model’s performance on this query is considered unstable, suggesting that the query is challenging and holds high learning potential. In this case, only the responses that exceed the verify_score threshold are retained to ensure data quality.
@@
      If the CV is less than or equal to 0.05, the query is deemed easy, and the model’s performance is stable. To maintain diversity and learnability within the dataset, we retain these queries based on their category. For other and multiturn data, we retain 50% of the samples at random to avoid excessive simplification of the dataset; for all other categories, such queries are discarded by default.
@@
Specifically, we maintain the same categories as defined in Stage I (math, code, science, instruction follow, and other), set a stringent verify_score threshold of 0.99, and apply a CV threshold of 0.05, retaining only those queries exceeding this variability threshold. Moreover, we exclusively select queries exhibiting the highest CV values, and for multiple model responses to a single query, we randomly sample only one response for training.
@@ 4.2 Annealing
We performed an annealing training experiment starting from the optimal 32B model obtained previously. Unlike Stage I, we adopted a more stringent data selection strategy, retaining only the highly challenging data characterized by a coefficient of variation (CV) greater than 0.05 (see Section 3.2). In this stage, we discontinued the use of the packing strategy to allow the model to focus exclusively on handling complex long-reasoning tasks for individual query-response pairs, while still maintaining a maximum sequence length of 32k tokens. We set a lower learning rate of $8\times 10^{-6}$ and trained the model for two epochs, further exploring the model’s potential on challenging long-reasoning tasks. Similarly, we employed a cosine learning rate scheduler with a linear warmup phase for the first 5% of the training steps, gradually decaying the learning rate to zero.

### v3

@@ 3.2 Data Selection
2:Category-specific verify_score threshold and fixed CV threshold ($CV_{t}$)
@@
11:     if CV $>$ $CV_{t}$ then $\triangleright$ Challenging query
@@
      If the CV exceeds the threshold of $CV_{t}$ [^4] Note that different category has different CV threshold., the model’s performance on this query is considered unstable, suggesting that the query is challenging and holds high learning potential. In this case, only the responses that exceed the verify_score threshold are retained to ensure data quality.
@@
      If the CV is less than or equal to $CV_{t}$, the query is deemed easy, and the model’s performance is stable. To maintain diversity and learnability within the dataset, we retain these queries based on their category. For other and multiturn data, we retain 50% of the samples at random to avoid excessive simplification of the dataset; for all other categories, such queries are discarded by default.
@@
Specifically, we maintain the same categories as defined in Stage I (math, code, science, instruction follow, and other), set a stringent verify_score threshold of 0.99, retaining only those queries exceeding $CV_{t}$. Moreover, we exclusively select queries exhibiting the highest CV values, and for multiple model responses to a single query, we randomly sample only one response for training.
@@ 4.2 Annealing
We performed an annealing training experiment starting from the optimal 32B model obtained previously. Unlike Stage I, we adopted a more stringent data selection strategy, retaining only the highly challenging data characterized by a coefficient of variation (CV) greater than $CV_{t}$ (see Section 3.2). In this stage, we discontinued the use of the packing strategy to allow the model to focus exclusively on handling complex long-reasoning tasks for individual query-response pairs, while still maintaining a maximum sequence length of 32k tokens. We set a lower learning rate of $8\times 10^{-6}$ and trained the model for two epochs, further exploring the model’s potential on challenging long-reasoning tasks. Similarly, we employed a cosine learning rate scheduler with a linear warmup phase for the first 5% of the training steps, gradually decaying the learning rate to zero.